gsmith2002:

What is the discriminant of the quadratic equation 0 = –x2 + 4x – 2? –4 8 12 24

Tranquility:

The discriminant of the equation ax^2 + bx + c is b^2-4ac If you recall correctly, that's the term under the square root in the quadratic formula

gsmith2002:

so -4 ?

gsmith2002:

hello

Tranquility:

Show your work

gsmith2002:

−b±b2−4(a)(c) √2(a)−4±42−4(−1)(−2) √2(−1)−4±8–√−2=−4±4×2 √−2=−4±22–√−2=−2±12–√−1=2±−12–√

Tranquility:

The discriminant is b^2-4ac It seems like you're trying to find the solutions to the equation

Tranquility:

@gsmith2002 You don't have the correct answer yet.